# One Demo Is All It Takes: Planning Domain Derivation with LLMs from A Single Demonstration

### v1

@@ A.8 Planning Time Limit Variation
Method Expert LLMTAMP LLMTAMP-FF LLMTAMP-FR RuleAsMem PDDLLM
Stack $95.5\pm 0.6$ $41.5\pm 4.3$ $56.8\pm 3.4$ $43.8\pm 3.8$ $84.3\pm 3.3$ $97.5\pm 1.6$
Unstack $87.3\pm 0.8$ $81.2\pm 1.1$ $85.0\pm 5.7$ $85.5\pm 5.4$ $79.8\pm 1.9$ $94.9\pm 0.5$
@@
Method Expert LLMTAMP LLMTAMP-FF LLMTAMP-FR RuleAsMem PDDLLM
Stack $97.7\pm 0.7$ $41.5\pm 4.3$ $76.7\pm 2.7$ $71.0\pm 2.7$ $85.5\pm 2.9$ $97.5\pm 1.6$
@@
Overall $94.4\pm 0.2$ $35.8\pm 0.4$ $55.6\pm 1.2$ $54.8\pm 0.7$ $69.9\pm 0.7$ $94.2\pm 0.5$
As outlined in the experimental design section, we set a default planning time limit of 50 seconds to reflect the real-time constraints commonly imposed on robotic systems. Although prior studies typically allow planning times on the order of minutes (Silver et al., 2020; Khodeir et al., 2023; Huang et al., 2024), the exact limits vary considerably. To assess the robustness of our approach under different time allowances, we evaluate performance at time limits of 25, 50, and 100 seconds. The results show that our method consistently outperforms all LLM-based baselines across all tested time settings and task types. Moreover, the planning domains derived by PDDLLM yield performance comparable to that of expert-designed domains under all time settings and across all tasks. All simulations were conducted on a system with an Intel Core i9-14900KF CPU without GPU acceleration, and all LLM prompting was performed using API services.
@@ A.9 Example of PDDLLM imagined predicates
A.10 Broader Impact
PDDLLM lowers the barrier to deploying intelligent robots by enabling automated planning domain generation from natural language, making task planning more accessible and adaptable across a wide range of real-world applications. This capability has the potential to accelerate robotics development by reducing reliance on expert knowledge and allowing non-specialists to teach robots through demonstrations and language, accelerating deployment in areas like manufacturing, home assistance, education, and healthcare. It also promotes inclusivity by enabling intuitive human-robot interaction, particularly for users without technical backgrounds. However, the method introduces risks related to safety, reliability, and interpretability, as automatically derived domains may misrepresent task requirements or lead to unintended behaviors, especially in high-stakes or safety-critical settings. These concerns highlight the need for human oversight, safety validation, and alignment mechanisms to ensure robust and trustworthy robot behavior. Overall, while PDDLLM advances the scalability and usability of autonomous systems, careful deployment practices are essential to mitigate potential societal risks.

### v3

@@ A.8 Planning Time Limit Variation
Method Expert LLMTAMP LLMTAMP-$F$ LLMTAMP-FR RuleAsMem PDDLLM
Stack $95.8\pm 0.2$ $41.5\pm 4.3$ $56.8\pm 3.4$ $43.8\pm 3.8$ $84.3\pm 3.3$ $97.5\pm 1.6$
Unstack $99.4\pm 0.6$ $81.2\pm 1.1$ $85.0\pm 5.7$ $85.5\pm 5.4$ $79.8\pm 1.9$ $94.9\pm 0.5$
@@
Method Expert LLMTAMP LLMTAMP-$F$ LLMTAMP-FR RuleAsMem PDDLLM
Stack $99.5\pm 0.5$ $41.5\pm 4.3$ $76.7\pm 2.7$ $71.0\pm 2.7$ $85.5\pm 2.9$ $97.5\pm 1.6$
@@
Overall $95.9\pm 0.1$ $35.8\pm 0.4$ $55.6\pm 1.2$ $54.8\pm 0.7$ $69.9\pm 0.7$ $94.2\pm 0.5$
As outlined in the experimental design section, we set a default planning time limit of 50 seconds to reflect the real-time constraints commonly imposed on robotic systems. Although prior studies typically allow planning times on the order of minutes (Silver et al., 2020; Khodeir et al., 2023; Huang et al., 2025), the exact limits vary considerably. To assess the robustness of our approach under different time allowances, we evaluate performance at time limits of 25, 50, and 100 seconds. The results show that our method consistently outperforms all LLM-based baselines across all tested time settings and task types. Moreover, the planning domains derived by PDDLLM yield performance comparable to that of expert-designed domains under all time settings and across all tasks. All simulations were conducted on a system with an Intel Core i9-14900KF CPU without GPU acceleration, and all LLM prompting was performed using API services.
@@ A.9 Example of PDDLLM imagined predicates
Table 10: Extended examples of imagined predicates for real-world domains.
Predicate Name
Predicates for Weight
Lift a box in warehouse (light-box ?a): 40 < weight < 70 kg
(heavy-box ?a): 70 < weight < 100 kg
Collect apple of different size (small-apple ?a): 100 < w < 166.7 g
(medium-apple ?a): 166.7 < w < 233.3 g
(large-apple ?a): 233.3 < w < 300 g
Predicates for Length
Fulfill a 15-cm coffee cup (needs-refill ?a): 0 < level < 3.75 cm
(half-full ?a): 3.75 < level < 7.5 cm
(nearly-full ?a): 7.5 < level < 11.25 cm
Measuring high school children height (medium-height ?a): 150 < h < 160 cm
(average-height ?a): 160 < h < 170 cm
(tall-height ?a): 170 < h < 180 cm
(very-tall ?a): 180 < h < 190 cm
Predicates for Air Quality
Plan when to travel (suitable-travel ?a): 0 < PM2.5 < 166.7
(moderate-air ?a): 166.7 < PM2.5 < 333.3
(high-risk ?a): 333.3 < PM2.5 < 500
Predicates for Brightness
Fix broken light in office (dimly-lit ?a): 0 < brightness < 150 lux
A.10 Test Problem Set Design
A.10.1 Dimensions of Difficulty
Task Diversity: To ensure robustness and broad applicability, we sampled over 1,200 planning tasks across nine distinct environments, including stacking, unstacking, rearrangement, alignment, parts assembly, color classification, burger cooking, and bridge building. The details of the tasks are provided in Section˜A.4. Each task category includes tasks of 3 to 20 objects, with 10 distinct tasks sampled for each object count. The resulting task plan lengths ranged from 6 to 510 steps, reflecting a wide spectrum of planning horizons. Our experiments spanned over 150 unique predicates.
Task Complexity: In task design, we consider two types of complexity: domain derivation complexity and planning complexity. Domain derivation complexity is determined by the number of predicates imagined and actions invented by PDDLLM; the more predicates and actions, the higher the complexity. Planning complexity is influenced by both the planning domain and the task. Given $n$ objects, a task plan of length $l$, and $m$ actions in the domain, the branching factor at each step is $m\times n$, resulting in a approximate complexity of $(m\times n)^{l}$. In Section˜A.10, we layout a table showcase the planning complexity (in approximate order of magnitude) of the most difficult problem in the category and the domain derivation complexity. The Tower of Hanoi task exhibits the greatest planning complexity, while bridge building presents the highest domain derivation complexity. In contrast, stack and unstack are simpler in both complexity measures.
Task
Stack
Unstack
Color
Classify
Alignment
Parts
Assembly
Rearrange
Burger
Cook
Bridge
Build
Tower
of Hanoi
Max
Planning
Complexity
$10^{58}$ $10^{58}$ $10^{71}$ $10^{58}$ $10^{58}$ $10^{137}$ $10^{48}$ $10^{36}$ $10^{307}$
Domain
Derivation
Complexity
90 90 111 94 92 96 114 128 94
A.10.2 Derivation of Bounded Domain Derivation Complexity
The predicate imagination complexity measures how many predicates need to be generated. Our method is generalizable to more complex cases for the following reasons:
Complexity calculation formula
The formula is:
$\text{complexity}=\sum_{i}n_{p,i}^{\,n_{\text{dim},i}}$
where $n_{\text{dim},i}$ is the number of dimensions included in predicate type $i$. For example, $n_{\text{dim},\text{color}}=3$ since color is represented by the three RGB channels. $n_{p,i}$ is the number of partitions along each dimension for predicate type $i$. The term $n_{p,i}^{\,n_{\text{dim},i}}$ gives the total number of predicates generated for type $i$, and summing over all predicate types yields the overall complexity of the predicate imagination process for the task.
The Method is Scalable as the Increase in Predicate Types Leads to Linear Growth
The number of predicate types grow linearly as the task complexity increases. The only exponential factor comes from the number of dimensions $n_{\text{dim}}$ per predicate type; however, in practice, this value is typically small (e.g., $\leq 3$).
High-Dimensional Features Can Be Decomposed to Reduce Complexity
For higher-dimensional features, as object pose $(x,y,z,\text{yaw},\text{pitch},\text{roll})$, we can decompose them into lower-dimensional predicate types (e.g., position and orientation). This reduces exponential complexity to additive linear terms, preserving scalability and enabling rich predicate representation.
Task Planning Does Not Require Many Partitions
Dividing features into too many partitions makes predicate implication nearly continuous, which defeats the purpose of symbolic abstraction. Here is an experiment:
  • 2 partitions: “cold” ($0<T<50^{\circ}C$), “hot” ($50<T<100^{\circ}C$)
  • 4 partitions: “cool” ($0<T<25^{\circ}C$), “warm” ($25<T<50^{\circ}C$), “hot” ($50<T<75^{\circ}C$), “near-boiling” ($75<T<100^{\circ}C$)
  • 100 partitions: …, “post-boiling” ($95<T<96^{\circ}C$), “stable-boil” ($96<T<97^{\circ}C$), …
For the same task “boil milk”, we compare the predicate imagination results by forcing the system to partition the feature space into 2, 4, and 100 intervals. It is evident that overly fine partitioning reduces interpretability and semantic efficiency, so the number of partitions $n_{p}$ per dimension is usually low.
A.11 Missing and Redundant Elements
A redundant predicate or action refers to one that, when removed from the planning domain, does not degrade planning performance. Conversely, a missing predicate or action is one whose absence significantly reduces the domain’s planning performance.
The reported percentages are computed by comparing the PDDLLM generated PDDL domains with the expert-designed PDDL domains. The percentage of missing predicates/actions reflects the completeness of the generated domain (i.e., whether all necessary components are present). The percentage of redundant predicates/actions reflects the optimality of the domain (i.e., whether unnecessary components have been excluded).
A.12 Real-world Demonstration Collection
To enable the demonstration collection in real-world, we build up a data collection system using Agilex Pika and ArUco markers. We collect both the end-effector trajectories of the human demonstrator and the pose of the target objects. These data are later mapped to symbolic predicates and actions in the PDDL domain, providing a bridge between raw demonstration data and high-level planning representations.
End-Effector Trajectory: The movement of the human operator’s end effector is recorded using the Agilex Pika Data Collection System. The system’s motion capture capability provides precise position and orientation measurements over time. These trajectories serve as a basis for identifying meaningful segments of motion that can be abstracted into candidate actions for the symbolic domain.
Object Properties and States: Physical characteristics of the objects, such as size, are obtained from the available CAD models, ensuring that the symbolic model is grounded in realistic object dynamics. In addition, the poses of the objects during demonstrations are captured using ArUco markers affixed to their surfaces. These pose estimates allow us to track state changes—such as whether an object has been grasped, moved, or placed—which are subsequently encoded as predicates in the PDDL formulation.